if sec theta=x+1/x then what is the value of sectheta+tantheta

Grade:11

1 Answers

Tanmay
6 years ago
\sec \theta =x+1/x \tan \theta =\sin\theta*\sec\theta =\sin\theta*x+1/x \tan\theta+\sec\theta=\sin\theta*x+1/x + x+1/x = (x+1)(\sin\theta+1)/x